### v1

# PaCE: Parsimonious Concept Engineering for Large Language Models

Abstract
Large Language Models (LLMs) are being used for a wide variety of tasks. While they are capable of generating human-like responses, they can also produce undesirable output including potentially harmful information, racist or sexist language, and hallucinations. Alignment methods are designed to reduce such undesirable output, via techniques such as fine-tuning, prompt engineering, and representation engineering. However, existing methods face several challenges: some require costly fine-tuning for every alignment task; some do not adequately remove undesirable concepts, failing alignment; some remove benign concepts, lowering the linguistic capabilities of LLMs. To address these issues, we propose Parsimonious Concept Engineering (PaCE), a novel activation engineering framework for alignment. First, to sufficiently model the concepts, we construct a large-scale concept dictionary in the activation space, in which each atom corresponds to a semantic concept. Given any alignment task, we instruct a concept partitioner to efficiently annotate the concepts as benign or undesirable. Then, at inference time, we decompose the LLM activations along the concept dictionary via sparse coding, to accurately represent the activation as a linear combination of the benign and undesirable components. By removing the latter ones from the activation, we reorient the behavior of LLMs towards alignment goals. We conduct experiments on tasks such as response detoxification, faithfulness enhancement, and sentiment revising, and show that PaCE achieves state-of-the-art alignment performance while maintaining linguistic capabilities. Our collected dataset for concept representations is available at https://github.com/peterljq/Parsimonious-Concept-Engineering.
1 Introduction
Large Language Models (LLMs) are useful for tasks as far ranging as question answering [76, 65], symbolic reasoning [57, 25], multi-modal synthesis [86, 41, 45], and medical diagnosis [85]. LLMs are typically pre-trained on a broad collection of textual corpora with the next-token prediction objective [55, 70], enabling them to generate human-like text. An important aspect of deploying pre-trained LLMs for real-world applications is preventing undesirable responses such as toxic language, hallucinations, and biased information through alignment methods, which aim to make AI systems behave in line with human intentions and values [28]. A common alignment approach is tuning LLMs with human feedback [56, 62] for better instruction-following capabilities. However, after such aligning, undesirable and harmful content can still be elicited from LLMs. For example, jailbreaking can produce hate speech and aggression [22, 32], stress-testing shows hallucinatory responses such as illogical statements [87], and various kinds of bias are not fully removed from LLM responses [19]. This emphasizes the need for further development towards aligned LLMs.
Overall, alignment methods can largely be categorized into: parameter fine-tuning, prompt engineering, and activation engineering. Parameter fine-tuning methods, such as low-rank adaptation [26] and knowledge editing [14, 73], involve updating the model parameters using datasets of input-response pairs [74]. Unfortunately, such computations over large datasets are often costly. Furthermore, whenever a new category of undesirable behaviors is identified or a new group of customers is acquired, the LLM supplier has to incur the cost of data creation and fine-tuning again. Prompt engineering attempts to manipulate the LLM’s reasoning with carefully designed instruction prompts [77, 81, 79]. However, effective instructions are commonly obtained through empirical trial-and-error, with no guarantee of coverage across tasks of different domains. Notably, recent works show that the instruction itself can be lengthy [39] or contain human errors [10, 61].
Figure 1: Our framework PaCE achieves alignment goals by sparse coding and adjusting vectors in the activation space of the LLM Decoder Layer (DCL).
Activation engineering, i.e., algorithms that modify the latent activations of LLMs, has emerged to alleviate high-cost and poor coverage of tasks. Recent work has shown that certain directions in the activation space of LLMs are associated with semantic concepts (c.f. §2.1). Thus, given an input prompt at inference time, modifying its neural activations towards or away from these directions controls the semantics of the model response. For example, methods based on Vector Addition (VecAdd) [67, 71, 88, 44, 54, 68, 69, 38] directly add multiples of a concept direction to a neural activation, while those based on Orthogonal Projection (OrthoProj) [88, 23] subtract from a neural activation its orthogonal projection onto a concept direction. Nonetheless, these methods face two major challenges. First, these methods inadequately model the geometry of the activation space, as we will detail in §2.2. Hence, they tend to either remove benign concepts, harming linguistic capability; or insufficiently remove undesirable concepts, thereby failing the alignment task. Second, for each alignment task, these methods typically only remove a single concept direction from the input activation vector, while there may be multiple concepts related to the alignment task.
To address these challenges, we propose Parsimonious Concept Engineering (PaCE), an activation engineering framework for alignment that i) enforces alignment goals effectively and efficiently, ii) retains linguistic capability, and iii) adapts to new alignment goals without costly parameter fine-tuning. PaCE consists of two stages: (1) Concept Construction and Partition, and (2) Activation Decomposition and Intervention (Figure 3). We summarize the procedure of PaCE below and highlight our contributions in bold.
  • Concept Dictionary Construction and Partition (§3.2): Since existing works only provide a limited number of concept directions, we collect a large concept dictionary, PaCE-1M, that consists of 40,000 concept directions extracted from over 1,200,000 context sentences. In particular, for each concept in the Brown Corpus [18], we use a knowledge-driven GPT [65, 45, 36] to propose contextual scenarios to describe the concept, and extract concept directions in the representation (activation) space [88] from the context sentences. This is done only once offline. Given an alignment task, we further instruct a GPT to automatically partition the concept directions in the dictionary into benign and undesirable directions.
  • Activation Decomposition and Intervention (§3.3): At inference time, given any user input prompt, we decompose the activations as a sparse linear combination of concept directions using sparse coding techniques. Notably, this allows for an efficient and accurate estimate of both undesirable and benign components in the activations, which is overlooked in previous activation engineering methods. By removing the undesirable components from the activations, we reorient the behavior of LLMs toward alignment goals, while maintaining their linguistic capability.
We evaluate PaCE on multiple alignment tasks including response detoxification, faithfulness enhancement, and sentiment revising (§4). We show that PaCE achieves state-of-the-art performance on these tasks, while retaining its linguistic capability at a comparable level. We further shed insights on the concept directions of PaCE-1M by showing that they are geometrically consistent with their concept semantics and a decomposition reveals the semantics of the activations.
2 Basics of Latent Space Engineering
As motivated above, in this paper we are interested in controlling LLMs by leveraging structures in their latent space. We begin by reviewing some basic properties of the latent space in §2.1. This lays the foundation for previous methods on latent space intervention in §2.2 as well as our method in §3.
Figure 2: To remove a concept direction ‘red’ from the latent code ‘red apple’ (left), prior works use i) orthogonal projection (middle right, (OrthoProj)), which may remove extra directions, or ii) vector addition (right, (VecAdd)), where it is hard to pick the edit strength $c$. Instead, PaCE explicitly models the concept dictionary in the latent space and use oblique projection (middle left).
2.1 The Latent Space and Its Linear Controllability
[…]
Remark 1 (${\mathcal{Z}}=$ Word Embeddings).
A classic setting where linear controllability shows up is that of word embeddings. Here, ${\mathcal{T}}$ is the vocabulary (say, the set of English words), ${\mathcal{Z}}$ contains some vectors in ${\mathbb{R}}^{d}$, and $g$ is a bijection between ${\mathcal{Z}}$ and ${\mathcal{T}}$. In the seminal work of Mikolov et al. [51], the authors observe that word embeddings learned by recurrent neural networks approximately enjoy relations such as ${\boldsymbol{z}}_{\text{king}}-{\boldsymbol{z}}_{\text{man}}+{\boldsymbol{z}}_{\text{woman}}\approx{\boldsymbol{z}}_{\text{queen}}$, where one can view ${\boldsymbol{z}}_{\text{woman}}-{\boldsymbol{z}}_{\text{man}}$ as the concept direction ${\boldsymbol{v}}\in{\mathcal{V}}$ and the control strength to be $c=1$. This observation is later extended to word embeddings of various networks and learning objectives such as word2vec [50], Skip-Grams [49, 35], GloVe [59], and Swivel [63]. On the theoretical front, a fruitful line of research has been devoted to understanding the emergence of such properties in word embeddings [3, 21, 2, 1, 17, 53].
Remark 2 (${\mathcal{Z}}=$ Neural Activations).
Modern neural architectures such as transformers have significantly boosted the linguistic performance of language models. Much of their success is attributed to the attention mechanism, which incorporates long-range context into the neural activations in transformers. This has motivated people to take ${\mathcal{Z}}$ as certain hidden states in transformers [^2] A variety of choices of layers have been explored in the literature; see, e.g., [67] for a comparison., and search for concept directions ${\mathcal{V}}$. This has led to a fascinating line of works supporting the empirical existence of ${\mathcal{V}}$: [6, 47] find directions that indicate truthful output, [68] finds directions for sentiments, [88] finds directions for emotions and honesty, and [54] finds directions for current player tile in a synthetic board game model. Interestingly, [58, 30] further offer theoretical models, under which the linear controllability shows up provably in the latent space of LLMs.
2.2 Controlling Language Models via Latent Space Engineering
The above findings have supported the development of practical methods to control the behavior of language models. As we will see, a key challenge there is to decide the correct control strength.
Vector Addition. The work of [67, 71, 88, 44, 54, 68, 38] proposes to add or subtract multiples of a concept direction from the latent code. For example, to remove hatred from ${\boldsymbol{z}}$, one performs
${\boldsymbol{z}}\mapsto{\boldsymbol{z}}-\hat{c}\cdot{\boldsymbol{v}}_{\text{hatred}},$ (VecAdd)
where $\hat{c}>0$ is a parameter of the strength of control. In principle, as the input prompt may contain a different ‘extent’ of the concept to be removed, $\hat{c}$ should depend on both the input prompt and the concept. Thus, in practice, one either tunes $\hat{c}$ per input prompt and concept, which is laborious, or one fixes a $\hat{c}$, which is sub-optimal. Indeed, this has been observed by the work [71]: In their Table 10, the optimal coefficients $\hat{c}$ are markedly different across the examples; see also their ‘discussion’ section.
Orthogonal Projection. The work of [5] proposed to remove gender bias in word embeddings by projecting the embeddings onto the orthogonal complement to a gender direction ${\boldsymbol{v}}_{\text{gender}}$:
${\boldsymbol{z}}\mapsto\Pi_{\operatorname{span}({\boldsymbol{v}}_{\text{gender}})^{\perp}}{\boldsymbol{z}}={\boldsymbol{z}}-\Pi_{\operatorname{span}(v_{\text{gender}})}{\boldsymbol{z}}.$ (OrthoProj)
Here, for any ${\boldsymbol{w}}\in{\mathbb{R}}^{d}$, $\operatorname{span}({\boldsymbol{w}})$ is the linear subspace spanned by ${\boldsymbol{w}}$, and for any linear subspace ${\mathcal{S}}\subset{\mathbb{R}}^{d}$, $\Pi_{{\mathcal{S}}}$ denotes the ortho-projector onto ${\mathcal{S}}$. Such an idea is later applied to neural activations of LLMs [23, 88]. Applying orthogonal projection to remove concept directions from latent codes may be reasonable: if directions corresponding to different concepts are orthogonal, then orthogonal projection does not remove directions from concepts other than the gender direction. That being said, there are often more concept directions presented, and they are not orthogonal. For example, [29] shows that causally related concepts only exhibit partial orthogonality for their directions.
To sum up, numerous attempts have been made to control the behavior of language models. However, existing methods either have a control strength parameter that is hard to tune or may remove extra concept directions. As we will see in the next section, these issues can be resolved by the proposed PaCE framework, which explicitly models the geometry of a large concept dictionary.
3 Our Method: Parsimonious Concept Engineering
Figure 3: Pipeline of PaCE has several major steps: Step 1 collects concept vectors and constructs the concept dictionary, Step 2 decomposes the activation vector of the given input by sparse coding to get concept coefficients, and Step 3 performs editing on the concepts towards reoriented response.
3.1 Activation Intervention via Overcomplete Oblique Projection
Can we efficiently remove one or more target concept directions from a given latent activation without affecting other concept directions present? To address this problem, our key insight is to model as many concept directions as possible, and then decompose the activation to estimate its components along these directions. Figure 2 presents an idealized visual example. Here, one is given a latent activation meaning ‘red apple’, and the goal is to remove the ‘red’ direction from the activation (left). As illustrated, orthogonal projection and vector addition tend to fail (middle right and right), as we discussed in §2.2. In contrast, by decomposing the activation along the concept directions of ‘red’ and ‘apple’, one can safely remove the component along ‘red’ without affecting that along ‘apple’ (middle left). This is related to the idea of oblique projection, which gives the name of this section.
That said, several challenges remain to be addressed. As motivated above, to accurately model semantic concepts, one needs to collect as many concept directions in the latent space as possible. Since existing works only provide a limited number of concept directions (as reviewed in Remark 2), we contribute by collecting a large dictionary of concept directions, which we will discuss in §3.2. Moreover, oblique projection is well-defined only when the concept directions are linearly independent, while concept directions are often dependent (as we show in §4.3) so the decomposition is not unique. §3.3 discusses our choice of decomposition algorithm to address this difficulty.
3.2 Knowledge-Driven Concept Dictionary
Concept Dictionary Construction. We take the top 40,000 words from the Brown Corpus [18] ranked by word frequency [4] as the concept collection $T$. For each concept $t_{i}\in T$, we prompt GPT-4 to generate around $30$ pieces of contextual stimuli $s_{i}=\{s_{i}^{1},s_{i}^{2},\cdots,s_{i}^{30},\cdots\}$ that are scenarios describing the concept. To enhance the diversity of the concept stimuli, we retrieve knowledge from Wikipedia [65, 45, 36] (as we detail in Appendix B.4) to augment the prompt of stimulus synthesis. Samples of concepts and their stimuli are shown in Figure 4 and Appendix Figure 11. For each concept $t_{i}$, we extract a direction ${\boldsymbol{v}}_{i}^{\ell}$ from the activations of its contextual stimuli at the $\ell$-th decoder layer of the LLM [88], which gives a dictionary ${\boldsymbol{D}}^{\ell}\in{\mathbb{R}}^{d\times n}$ per layer (detailed in Appendix B.2).
Task-Driven Dictionary Partition. Given an alignment task, we further instruct GPT-4 as a concept partitioner to classify whether a concept needs to be removed from the input representation. To take detoxification as an example, the concept ‘harmful’ is highly correlated to the toxic response (hence needs removal) while benign concepts such ‘bird’ and ‘laptop’ will remain. That is, the instructed GPT-4 partitions the concepts into undesirable and benign to the alignment tasks. The full prompting templates of concept synthesis and partitioning are shown in Appendix E. In the next sub-section, we describe the notations and usages of the annotated concept dictionary.
Figure 4: Examples of the constructed concepts and their partition for the detoxification task sampled from our PaCE-1M.
3.3 Overcomplete Oblique Projection via Sparse Coding
Now that we have a dictionary ${\boldsymbol{D}}=[{\boldsymbol{v}}_{1},\dots,{\boldsymbol{v}}_{n}]\in{\mathbb{R}}^{d\times n}$ of $n$ concepts directions [^3] For notational simplicity, we discuss sparse coding for a single ${\boldsymbol{D}}$; Algorithm 2 deals with multiple layers., where each ${\boldsymbol{v}}_{i}$ is a concept direction of known semantic meaning. Given a latent activation ${\boldsymbol{z}}^{\text{in}}$ coming from the user input, how can we control it via oblique projection?
Oblique Projection. The general paradigm of oblique projection can be stated as follows.
  • Step $1$-Decomposition: Find $c_{1}^{\text{in}},\dots,c_{n}^{\text{in}}\in{\mathbb{R}}$ such that ${\boldsymbol{z}}^{\text{in}}=c_{1}^{\text{in}}{\boldsymbol{v}}_{1}+\dots+c_{n}^{\text{in}}{\boldsymbol{v}}_{n}+{\boldsymbol{r}}^{\text{in}}$ by solving
    ${\boldsymbol{c}}^{\text{in}}\in\mathop{\rm argmin}_{\boldsymbol{c}}\frac{1}{2}\|{\boldsymbol{z}}^{\text{in}}-{\boldsymbol{D}}{\boldsymbol{c}}\|_{2}^{2}+\Omega({\boldsymbol{c}}),$ (1)
    where $\Omega({\boldsymbol{c}})$ is a sparsity-promoting regularizer that we will discuss soon. Then, each coefficient $c_{i}^{\text{in}}$ for $i\in\{1,\dots,n\}$ can be viewed as how much the concept represented by ${\boldsymbol{v}}_{i}$ is in ${\boldsymbol{z}}^{\text{in}}$, and ${\boldsymbol{r}}^{\text{in}}$ is the residual that is not explained by ${\boldsymbol{D}}$.
  • Step 2-Intervention: Obtain the controlled coefficients $c_{1}^{\text{ctrl}},\dots,c_{n}^{\text{ctrl}}\in{\mathbb{R}}$, where $c_{i}^{\text{ctrl}}$ is set to $c_{i}^{\text{in}}$ if the concept of ${\boldsymbol{v}}_{i}$ is benign to the control task and 00 if undesirable (which has been decided offline in §3.2). Then, synthesize a new latent code using the modified coefficients and the residual by taking ${\boldsymbol{z}}^{\text{ctrl}}=c_{1}^{\text{ctrl}}{\boldsymbol{v}}_{1}+\dots+c_{n}^{\text{ctrl}}{\boldsymbol{v}}_{n}+{\boldsymbol{r}}^{\text{in}}$.
The synthesized ${\boldsymbol{z}}^{\text{ctrl}}$ will replace ${\boldsymbol{z}}^{\text{in}}$ to be passed on to the next layer of the neural network.
Remark 3 ((OrthoProj, VecAdd) $=$ Special Cases of Oblique Projection).
If one restricts ${\boldsymbol{D}}$ to contain only the undesirable concept directions (i.e., the ones to be removed from the latent code), and further takes $\Omega(\cdot)$ to be a constant function, it can be shown that oblique projection reduces to the special case of orthogonal projection (OrthoProj). On the other hand, if ${\boldsymbol{D}}$ contains only one undesirable concept direction, and $\Omega(\cdot)$ is $\lambda\|\cdot\|_{2}^{2}$ for some regularization strength $\lambda\in{\mathbb{R}}$, then oblique projection recovers vector addition (VecAdd), by setting $\lambda$ equal to $\hat{c}$ in (VecAdd). We provide proofs in Section B.1. As we will see next, our method differs from these two in having a larger dictionary and a sparsity-promoting regularizer.
Overcomplete Oblique Projection. As mentioned in §3.1, when the concept directions are linearly independent, then there is a unique decomposition of the latent code along the concept directions. However, often the concept directions can be dependent or nearly so, leading to infinitely many decompositions or numerical issues. To address this issue, we leverage the idea of sparse coding: natural signals are typically generated from sparse linear combinations of dictionary atoms, and pursuing a sparse decomposition reveals certain aspects of the underlying signal despite the dictionary being overcomplete (i.e., the system is underdetermined) [^4] For example, identifying which atoms or which blocks of atoms that the underlying signal is from [16].. This has been explored in a fruitful line of research in machine learning and computer vision (see textbooks [13, 72, 78] and references therein). Following this idea, we solve (1) with the regularizer $\Omega({\boldsymbol{c}})$ chosen to be the elastic net, i.e.,
$\Omega({\boldsymbol{c}})=\alpha\left\lparen\tau\|{\boldsymbol{c}}\|_{1}+(1-\tau)\frac{1}{2}\|{\boldsymbol{c}}\|_{2}^{2}\right\rparen,$ (2)
where $\tau\in[0,1]$ and $\alpha>0$ are parameters that control the sparsity of the solution. This problem is efficiently solved via an active-set algorithm that leverages the sparsity of the solution [82]. Pursuing sparse codes that emerges from the data is often known as parsimonious representation learning [42], which gives rise to the name PaCE of our overall framework. We summarize the online intervention process in Algorithms 1 and 2, and the overall PaCE procedure in Algorithm 3 in the Appendix.
Input: Latent vector ${\boldsymbol{z}}^{\text{in}}$, dictionary ${\boldsymbol{D}}$, index set $I$ of undesirable concepts
[…]
  $\triangleright$ Analysis
${\boldsymbol{r}}^{\text{in}}={\boldsymbol{z}}^{\text{in}}-{\boldsymbol{D}}{\boldsymbol{c}}^{\text{in}}$
  $\triangleright$ Residual
${\boldsymbol{c}}^{\text{ctrl}}=\Pi_{\langle e_{i},\forall i\in I\rangle^{\perp}}{\boldsymbol{c}}^{\text{in}}$
  $\triangleright$ Control
${\boldsymbol{z}}^{\text{ctrl}}={\boldsymbol{r}}^{\text{in}}+{\boldsymbol{D}}{\boldsymbol{c}}^{\text{ctrl}}$
  $\triangleright$ Synthesis
return Intervened latent vector ${\boldsymbol{z}}^{\text{ctrl}}$
Algorithm 1 Overcomplete Oblique Projection (ObliqProj)
Input: Pre-trained LLM with $L$ decoder layers (DCL) to decompose, input tokens ${\boldsymbol{E}}$, dictionaries $\{{\boldsymbol{D}}^{\ell}\}_{\ell=1}^{L}$, index set $I$ of undesirable concepts
${\boldsymbol{z}}_{1}=\text{LayersBeforeDCL}({\boldsymbol{E}})$
For $\ell\leftarrow 1,2,\dots,L$:
${\boldsymbol{z}}^{\ell}=\text{ObliqProj}({\boldsymbol{z}}^{\ell},{\boldsymbol{D}}^{\ell},I)$
  $\triangleright$ Algorithm 1
${\boldsymbol{z}}^{\ell+1}=\text{DCL}^{\ell}({\boldsymbol{z}}^{\ell})$
$\indent{\boldsymbol{e}}=\text{LayersAfterDCL}({\boldsymbol{z}}^{L+1})$
return Output token ${\boldsymbol{e}}$
Algorithm 2 PaCE Activation Intervention
4 Experimental Results
We evaluate the effectiveness of PaCE on downstream tasks including Detoxification, Faithfulness Enhancement, and Sentiment Refinement. We then analyze the sampled activation space, enabled by our large collection of concept vectors. We provide implementation details in Appendix B.4.
4.1 Improving Safety by Response Detoxification
Figure 5: An example of jailbreaking LLaMA2-7B-Chat and detoxification by PaCE. PaCE successfully detoxifies the response while maintaining the instruction-following capability.
Here we perform activation manipulation using our framework PaCE for detoxifying LLM responses. An example of our detoxification is shown in Figure 5: LLaMA2-7B-Chat is prompted with the malicious intent (i.e., jailbreaking) and parts of the response of the vanilla LLM (vanilla response) are generally considered manipulative and ill-intent. Our PaCE response pivots from a harmful to a harmless style and makes harmless suggestions. Appendix D.1 shows additional concrete examples.
Setup. For baselines, Prompting directly instructs LLM not to output sentences relevant to the list of top undesirable concepts (template in Appendix B), VecAdd subtracts the concept vector ‘harmful’ from the activation of the input, and OrthoProj performs projection on the orthogonal complement of the concept vector ‘harmful’. Note that, if we directly apply OrthoProj and VecAdd over the large collection of top undesirable concepts (e.g., 50 concepts) with no decomposition analysis, the input representation will significantly diverge from the original ones since every activation vector is of a similar scale, and the LLM’s linguistic capabilities will degrade. We compare our method in defending maliciousness against activation manipulation methods (§2.2) on the SafeEdit [73] dataset with its safety scorer. For every response, the benchmark’s safety scorer rates between 00 and $1$ (higher is safer). We use the effective set where the original safety score is lower than $50\%$ (i.e., the successful attacks if binarily classified).
Safety Responses. The evaluation has nine categories: Political Sensitivity (PS), Pornography (PG), Ethics and Morality (EM), Illegal Activities (IA), Mental Harm (MH), Offensiveness (OF), Physical Harm (PH), Privacy and Property (PP), and Unfairness & Bias (UB). As shown in Table 1, for LLaMa2-7B, PaCE improves by 60-80% over the vanilla method in categories including IA, MH, OF, PH, PP, and UB. When compared to other methods, PaCE performs competitively and improves by 6-20%. While our method did not perform the best in PS, PG, and EM, the gap for those categories is relatively small considering the significant overall gains. Notably, for LLaMA2-13B which has more parameters and a presumably more structured latent space, PaCE dominates other methods in all categories, demonstrating the necessity for respecting the latent structures when modifying representations. Finally, Table 3 shows the contribution of design choices in PaCE, and Table 3 shows the effect of the dictionary size on the performance. We observe clear improvement after each design choice is progressively added to PaCE. Appendix B.5 includes the details of these ablation studies.
Linguistic Capability. To validate that the detoxified representations of PaCE are still effective on general linguistic capability, we also evaluate the responses by N-gram fluency and perplexity. Furthermore, we apply PaCE to detoxify MMLU questions (which are naturally unharmful) to show that the detoxification will not significantly degrade the LLM’s reasoning capability. We observe that the MMLU response accuracy of PaCE is the highest among all activation manipulation baselines.
Figure 6: The detoxification performances for LLaMA2-13B w.r.t. the dictionary size.
Efficiency. Table 3 shows that PaCE is more time-efficient compared to the OrthoProj which also projects the concept vector onto the input vector. PaCE sees a three times speed improvement in average time per response and a two times improvement over average time per word when compared to OrthoProj. While PaCE is computationally slower than VecAdd, we argue the performance gain in a majority of the categories is a benefit that outweighs this particular shortcoming.
Table 1: Detoxification evaluation for PaCE, representation manipulation, and training-free baselines. The best performance of each category is in bold and the second best is underlined.
Target Model Method Safety (%, $\uparrow$) Linguistic Capability
PS PG EM IA MH OF PH PP UB Fluency ($\uparrow$) Perplexity ($\downarrow$) MMLU (%, $\uparrow$)
LLaMA-7B-Chat Vanilla [70] 17.6 19.5 10.1 7.79 11.3 17.2 22.6 11.8 17.2 7.70 3.51 43.4
Prompting [70] 82.5 47.3 57.8 65.2 75.1 54.8 72.0 72.4 56.1 7.50 3.04 15.4
VecAdd [67, 71, 88] 50.9 58.9 59.0 53.9 66.1 55.0 60.7 61.7 66.4 6.58 7.58 29.0
OrthoProj [23, 88] 50.7 57.9 50.2 47.5 67.0 50.1 74.9 65.7 66.4 7.46 3.73 34.1
PaCE (Ours) 69.6 46.2 58.2 75.3 94.2 62.3 80.8 72.8 88.3 8.07 3.52 37.1
LLaMA2-13B-Chat Vanilla [70] 8.01 23.7 13.6 19.8 18.3 21.6 13.6 14.0 16.7 7.66 2.48 54.9
Prompting [70] 35.8 68.3 59.3 52.5 73.5 23.4 78.0 71.1 66.5 7.63 2.22 52.1
VecAdd [67, 71, 88] 76.6 71.4 70.0 64.3 87.2 66.9 47.4 74.5 71.1 7.46 2.75 51.6
OrthoProj [23, 88] 51.1 82.6 50.6 72.4 52.3 58.0 51.4 65.1 75.5 7.29 2.88 52.9
PaCE (Ours) 93.7 97.9 97.7 94.9 98.9 96.6 99.3 90.8 98.9 7.52 2.85 54.1
Table 2: Computation time (in seconds) evaluation for PaCE and representation manipulation baselines. We observe that, compared to OrthoProj which also projects the concept, our PaCE is more time-efficient for trustworthiness control.
Method LLaMA2-7B-Chat LLaMA2-13B-Chat
Time per Response Time per Token Time per Response Time per Token
Vanilla 12.4 0.041 20.7 0.076
VecAdd 16.3 0.062 29.1 0.109
OrthoProj 143.7 0.514 221.6 0.780
PaCE (Ours) 44.8 0.119 50.3 0.149
Table 3: Ablation study for PaCE on the detoxifying LLaMA2-7B. Starting from a small emotion dictionary and manually selected concepts for removal, each subsequent design leads to better performance.
Method Safety (%, $\uparrow$) Fluency ($\uparrow$)
PaCE (LLaMA2-7B-Chat) 50.2 7.26
+ Decomposition on $10^{4}$ Concepts 57.6 7.58
+ Clustering of Concepts 62.3 7.63
+ Concept Partitioner 65.1 7.70
+ Removal of Top $50$ Concepts 76.5 8.07
4.2 Improving Faithfulness and Removing Negative
Sentiment
We evaluate the framework based on the response’s faithfulness and sentiment when input prompts requests for information involving biographical facts or minority social groups. Faithfulness reflects the level of factuality in the generation, and sentiment describes the emotional tone behind the generation. In short, we find PaCE effective in improving the faithfulness and removing negative sentiment in LLMs’ outputs. We describe the setup, metrics and method below.
Setup. Faithfulness: We use the FactScore suite and the fact evaluator for faithful biography generation [52]. The suite is divided into labeled and unlabeled subsets used in different sections of the original paper. Our table reports the Labeled Score (LS), the total number of Labeled Atomic Facts (LAF), Unlabeled Score (US), and the total number of unlabeled Atomic Facts (LAF). Sentiment: We use the HolisticBias suite [66] and hate speech evaluator [64] to measure the sentiment of the response to underrepresented descriptors. The reported numbers are the average of non-negative sentiment scores for underrepresented groups categorized by Gender (GN), Occupation (OC), and Nationality (NT). During the sentiment revising, the concept setups for all approaches follow the detoxification setup. For the faithfulness experiments, PaCE removes the top 50 undesirable (hallucinatory) concepts ranked by the GPT partitioner. The Prompting approach instructs the LLM not to output sentences relevant to these top concepts. The VecAdd and OrthoProj approaches operate on the concept vector of ‘fabrication’.
Figure 7: Concept coefficients solved by PaCE are an interpretable interface, and they are further used for activation intervention.
Results. Our results are shown in Table 4. For both 7B and 13B models, PaCE achieves more factual responses and improves the sentiment according to most metrics. For linguistic performance, our method ranks right after the Vanilla method for the larger 13B model, and achieves comparable results for LLaMA2-7B. Overall, we argue PaCE is an effective method for improving faithfulness and sentiment revising.
Table 4: Faithfulness and Fairness evaluation for PaCE, representation manipulation, and training-free baselines. The best performance of each category is in bold and the second best is underlined.
Target Model Method Fact ($\uparrow$) Sentiment (%, $\uparrow$) Linguistic Capability
LS (%) LAF US (%) UAF GN OC NT Fluency ($\uparrow$) Perplexity ($\downarrow$) MMLU (%, $\uparrow$)
LLaMA2-7B-Chat Vanilla [70] 18.4 45.1 15.4 37.4 51.5 69.2 56.4 7.20 2.49 43.4
Prompting [70] 28.6 40.6 20.4 49.0 53.1 62.3 56.6 7.25 2.87 16.3
VecAdd [67, 71, 88] 16.2 46.1 10.3 52.2 55.2 68.5 58.3 7.09 3.91 30.6
OrthoProj [23, 88] 21.9 49.7 26.2 45.9 54.9 75.1 60.1 7.21 2.76 34.1
PaCE (Ours) 27.7 65.9 30.8 73.3 66.2 79.7 69.9 7.91 2.88 38.4
LLaMA2-13B-Chat Vanilla [70] 44.1 39.6 41.8 38.5 50.2 70.3 58.1 7.63 2.41 54.9
Prompting [70] 61.6 24.5 47.5 20.0 46.1 73.8 59.4 7.46 2.45 52.4
VecAdd [67, 71, 88] 24.5 49.2 14.9 68.9 56.2 72.9 58.7 6.92 2.78 50.9
OrthoProj [23, 88] 59.3 52.8 43.2 51.7 57.7 75.1 63.3 7.26 2.66 51.1
PaCE (Ours) 64.8 53.0 76.4 55.1 63.4 76.5 67.5 7.48 2.43 53.1
Figure 8: The Representation (Activation) Space of LLaMA2-13B-Chat with the first 10000 Concepts from PaCE-1M. Appendix Figure 15 shows the zoom-in version. The visualization is the first two dimensions of UMAP of the concept vectors. We observe that concepts of similar semantics are clustered together, indicating that the activation space has semantic structures.
4.3 Representation Space Sampled by PaCE-1M
Our collected dataset of conceptual representations enables us to investigate the geometry and potential applications of the representation (activation) space.
Interpretability. Concept bottlenecks, where the model input is mapped to a list of human-readable concepts for interpretable decisions, are widely adopted for intervening model behaviors [34, 80, 7]. PaCE’s decomposition of the input prompt onto the large-scale concept dictionary also enables us to investigate the LLM’s internal behavior regarding the input prompt. Figure 7 shows the PaCE-solved weights for top nice concepts (in the order of absolute magnitude) in the activation space for an input prompt. The decomposition indicates that the target LLM corresponds the prompt toward the concept ‘irrational’ and against ‘conclusive’, which enables PaCE to execute the followed-up activation intervention (e.g., remove the concept ‘irrational’ by §3.3).
Concept Clustering and Retrieval. Here we explore the semantic structure of the activation space of the LLaMA2-13B-Chat by visualizing the first 10,000 concepts from the PaCE-1M dataset. We apply a dimensionality reduction method UMAP [48] on the concept vectors and visualize the first two dimensions in Figure 8. Concept vectors with similar semantics appear to be close to each other: e.g., in Figure 8 (1), concepts such as ‘college’, ‘university’, ‘Academy’, and ‘Institute’ are related to Education and they are close in the UMAP space. Notably, concepts of different semantics are clearly separated: those related to Education, Countries/States, Cities, Food and Clothing, and Positive Emotions respectively form distinct clusters. In particular, while concepts relevant to geography are closely clustered in Figure 8 (2), we observe a clear boundary between concepts related to Countries/States and those to Cities. These semantic structures indicate that the activation space sampled by our PaCE-1M dataset can capture and organize semantic information of the concepts, enabling further analysis and manipulations in PaCE. Figure 10 further reports the concept retrieval by evaluating the distance between a target concept with other concept vectors in the activation space. We observe organizational structure from the concept clusters based on their semantics. For instance, vectors for the concept ‘affection’ and ‘friendship’, are geometrically close and semantically relevant to the concept ‘love.’ Zooming out, such semantic structures are observed throughout the activation spaces of LLaMA2, and we conjecture they generalize to those in other LLMs. We provide more details of clustering and retrieval in Appendix C.2 and Appendix C.3.
Figure 9: Number of tokens per response across different intervention methods and LLM models.
Figure 10: The top 10 retrieved concepts using the similarity score in the sampled activation space. We observe close coherence between the target concept and retrieved concepts.
5 Conclusion and Discussion
In this paper, we present PaCE, an activation engineering framework designed for aligning LLMs by effectively and efficiently addressing undesirable representations while retaining linguistic capabilities. By constructing a large-scale concept dictionary and leveraging sparse coding for activation decomposition, PaCE opens up new research avenues for training-free LLM alignment. Our experiments on tasks such as response detoxification, faithfulness enhancement, and sentiment revising demonstrate that PaCE achieves state-of-the-art performance compared to existing representation manipulation approaches. PaCE not only ensures alignment with less cost but also adapts to evolving alignment goals without significantly compromising the LLM’s linguistic proficiency. We open-source the PaCE-1M dataset to facilitate future research and practical applications of LLM alignment, and will release the source code soon. We further elaborate on the potential limitations, societal impacts, and future works of PaCE in Appendix B.6.
Acknowledgments and Disclosure of Funding
This research was supported by ARO MURI contract W911NF-17-1-0304, DARPA GARD HR001119S0026, DARPA RED HR00112090132, ODNI IARPA HIATUS Program Contract #2022-22072200005, and NSF Graduate Research Fellowship #DGE2139757. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of NSF, ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Supplementary Material
Appendix A Structure of The Appendix
The appendix is structured as follows:
Appendix B describes details of our PaCE framework, including proofs of propositions and a comprehensive explanation of the framework’s algorithm.
Appendix C elaborates on the PaCE-1M dataset, demonstrating the structure of the dataset with explorations of subspace clustering to analyze the dataset.
Appendix D presents textual results, including visualizations of baseline comparisons and samples of concept clusters.
Appendix E shows the instruction templates used for GPT-4 to synthesize and partition concepts.
Appendix B Details of PaCE Framework
This section validates the propositions of the PaCE framework discussed in §3.3, followed by descriptions of how to extract representations and the algorithm of the whole procedures of PaCE.
B.1 Proofs of Oblique Projection Recovers Vector Addition and Orthogonal Projection
Proposition 1.
Let ${\boldsymbol{D}}\in{\mathbb{R}}^{d\times n}$ be a dictionary matrix and ${\boldsymbol{z}}\in{\mathbb{R}}^{d}$ a latent code. Then, any solution ${\boldsymbol{c}}^{*}$ of the optimization problem
$\min_{\boldsymbol{c}}\left\lVert{\boldsymbol{z}}-{\boldsymbol{D}}{\boldsymbol{c}}\right\rVert_{2}^{2}$ (3)
satisfies ${\boldsymbol{D}}{\boldsymbol{c}}^{*}=\Pi_{\operatorname{range}({\boldsymbol{D}})}{\boldsymbol{z}}$. Therefore, the map ${\boldsymbol{z}}\mapsto{\boldsymbol{z}}-{\boldsymbol{D}}{\boldsymbol{c}}^{*}({\boldsymbol{z}})$ is the same as ${\boldsymbol{z}}\mapsto{\boldsymbol{z}}-\Pi_{\operatorname{range}({\boldsymbol{D}})}{\boldsymbol{z}}={\boldsymbol{z}}\mapsto\Pi_{\operatorname{range}({\boldsymbol{D}})^{\perp}{\boldsymbol{z}}}$ in (OrthoProj).
Proof.
Note that ${\boldsymbol{I}}=\Pi_{\operatorname{range}({\boldsymbol{D}})}+\Pi_{\operatorname{range}({\boldsymbol{D}})^{\perp}}$. Therefore, the objective of (3) can be written as
$\left\lVert{\boldsymbol{z}}-{\boldsymbol{D}}{\boldsymbol{c}}\right\rVert_{2}^{2}$ $=\left\lVert\Pi_{\operatorname{range}({\boldsymbol{D}})^{\perp}}{\boldsymbol{z}}+\Pi_{\operatorname{range}({\boldsymbol{D}})}{\boldsymbol{z}}-{\boldsymbol{D}}{\boldsymbol{c}}\right\rVert_{2}^{2}$
$=\left\lVert\Pi_{\operatorname{range}({\boldsymbol{D}})^{\perp}}{\boldsymbol{z}}\right\rVert_{2}^{2}+\left\lVert\Pi_{\operatorname{range}({\boldsymbol{D}})}{\boldsymbol{z}}-{\boldsymbol{D}}{\boldsymbol{c}}\right\rVert_{2}^{2}+2\langle\Pi_{\operatorname{range}({\boldsymbol{D}})^{\perp}}{\boldsymbol{z}},\Pi_{\operatorname{range}({\boldsymbol{D}})}{\boldsymbol{z}}-{\boldsymbol{D}}{\boldsymbol{c}}\rangle,$
[…]
$\langle\Pi_{\operatorname{range}({\boldsymbol{D}})^{\perp}}{\boldsymbol{z}},\Pi_{\operatorname{range}({\boldsymbol{D}})}{\boldsymbol{z}}-{\boldsymbol{D}}{\boldsymbol{c}}\rangle=\langle{\boldsymbol{z}},\Pi_{\operatorname{range}({\boldsymbol{D}})^{\perp}}\left\lparen\Pi_{\operatorname{range}({\boldsymbol{D}})}{\boldsymbol{z}}-{\boldsymbol{D}}{\boldsymbol{c}}\right\rparen\rangle=0.$
Therefore, problem (3) is equivalent to optimizing
$\left\lVert\Pi_{\operatorname{range}({\boldsymbol{D}})}{\boldsymbol{z}}-{\boldsymbol{D}}{\boldsymbol{c}}\right\rVert_{2}^{2},$
which is lower bounded by 00. This lower bound is realizable since $\Pi_{\operatorname{range}({\boldsymbol{D}})}{\boldsymbol{z}}\in\operatorname{range}({\boldsymbol{D}})$. Thus, any minimizer ${\boldsymbol{c}}^{*}$ must realize this lower bound, meaning $\Pi_{\operatorname{range}({\boldsymbol{D}})}{\boldsymbol{z}}={\boldsymbol{D}}{\boldsymbol{c}}$. So we are done. ∎
Proposition 2.
Let ${\boldsymbol{D}}$ contain only one concept direction ${\boldsymbol{v}}\in{\mathbb{R}}^{d}$. Let ${\boldsymbol{z}}\in{\mathbb{R}}^{d}$ be a latent code, and $\lambda>-1$ a regularization strength. Then, the solution $c^{*}\in{\mathbb{R}}$ of the optimization problem
$\min_{{\boldsymbol{c}}}\left\lVert{\boldsymbol{z}}-{\boldsymbol{D}}{\boldsymbol{c}}\right\rVert_{2}^{2}+\lambda\left\lVert{\boldsymbol{c}}\right\rVert_{2}^{2}\ \ \Leftrightarrow\ \ \min_{c}\left\lVert{\boldsymbol{z}}-c{\boldsymbol{v}}\right\rVert_{2}^{2}+\lambda c^{2}$ (4)
is given by $c^{*}=\frac{\langle{\boldsymbol{z}},{\boldsymbol{v}}\rangle}{\lambda+1}$. Therefore, the map ${\boldsymbol{z}}\mapsto{\boldsymbol{z}}-{\boldsymbol{D}}{\boldsymbol{c}}^{*}({\boldsymbol{z}})$ recovers (VecAdd): the former is the same as ${\boldsymbol{z}}\mapsto{\boldsymbol{z}}-\eta_{\lambda}{\boldsymbol{v}}_{+}$, where one can set any $\eta_{\lambda}>0$ by properly choosing $\lambda>-1$, and ${\boldsymbol{v}}_{+}$ is defined as ${\boldsymbol{v}}$ if $\langle{\boldsymbol{v}},{\boldsymbol{z}}\rangle>0$ and $-{\boldsymbol{v}}$ otherwise.
Proof.
Note that the objective of (4) is simply a univariate quadratic function of $c$:
$\|{\boldsymbol{z}}\|_{2}^{2}-2\langle{\boldsymbol{z}},{\boldsymbol{v}}\rangle c+(\lambda+1)c^{2}.$
This has a unique minimizer $c^{*}=\frac{\langle{\boldsymbol{z}},{\boldsymbol{v}}\rangle}{\lambda+1}$ since $\lambda+1>0$ by assumption. To prove the second part of the proposition, note that
${\boldsymbol{z}}-{\boldsymbol{D}}{\boldsymbol{c}}^{*}({\boldsymbol{z}})={\boldsymbol{z}}-c^{*}({\boldsymbol{z}}){\boldsymbol{v}}={\boldsymbol{z}}-\frac{\langle{\boldsymbol{z}},{\boldsymbol{v}}\rangle}{\lambda+1}{\boldsymbol{v}}={\boldsymbol{z}}-\frac{\left\lvert\langle{\boldsymbol{z}},{\boldsymbol{v}}\rangle\right\rvert}{\lambda+1}\cdot\left\lparen{\boldsymbol{v}}\operatorname{sign}\left\lparen\langle{\boldsymbol{z}},{\boldsymbol{v}}\rangle\right\rparen\right\rparen.$ ) ) . (5)
Define $\eta_{\lambda}:=\frac{\left\lvert\langle{\boldsymbol{z}},{\boldsymbol{v}}\rangle\right\rvert}{\lambda+1}$ and ${\boldsymbol{v}}_{+}:={\boldsymbol{v}}\operatorname{sign}(\langle{\boldsymbol{z}},{\boldsymbol{v}}\rangle)$. One can see that by varying $\lambda\in(-1,+\infty)$, $\eta_{\lambda}$ can take any value in $(0,\infty)$. This concludes the proof. ∎
B.2 Extracting Concept Directions and Constructing Dictionary
Recall from §3.2 that for each concept $t_{i}$, we have collected a set of context stimuli (i.e., sentences that describe $t_{i}$) $s_{i}=\{s_{i}^{0},s_{i}^{1},\cdots,s_{i}^{N_{s}}\}$. This totals $40,000$ concepts and more than $1,200,000$ context stimuli.
To obtain a vector for each concept, we follow the representation reading algorithm [88] to map the concept to the hidden states of LLM decoder layers. We describe the algorithm here for completeness. Each context sentence $s_{i}^{j}$ together with the concept $t_{i}$ is first plugged into a pre-defined prompt template, producing $\bar{s}_{i}^{j}$.
Consider the <concept $t_{i}$> in the following scenario: Scenario: <stimulus $s_{j}^{j}$> Answer:
For any prompt $p$, denote by $f^{\ell}(p)$ the activation of the last token at the $l$-th layer of the LLM when the input is $p$. Then, to extract a vector for concept $t_{i}$, one looks at the activations of pairs of stimuli
$X_{i}^{\ell}:=\left\{\Pi_{{\mathbb{S}}^{d-1}}\left\lparen f^{\ell}(\bar{s}_{i}^{j})-f^{\ell}(\bar{s}_{i^{\prime}}^{j^{\prime}})\right\rparen:\forall i^{\prime}\neq i,\quad\forall j,j^{\prime}\right\},$ (6)
where $\Pi_{{\mathbb{S}}^{d-1}}(\cdot)$ is the projection onto the unit sphere, used to normalize the difference vectors. In practice, the work [88] uses a downsampled subset of $X_{i}^{\ell}$ rather than the entire $X_{i}^{\ell}$. We obtain the direction ${\boldsymbol{v}}_{i}^{\ell}$ of concept $i$ at layer $\ell$ by applying PCA on the set $X_{i}^{\ell}$, and taking the first principal direction; note that $\left\lVert{\boldsymbol{v}}_{i}^{\ell}\right\rVert_{2}=1$. Then, we construct the dictionary ${\boldsymbol{D}}^{\ell}=[{\boldsymbol{v}}_{1}^{\ell},\dots,{\boldsymbol{v}}_{n}^{\ell}]\in{\mathbb{R}}^{d\times n}$ of layer $\ell$, and doing this for all layers gives $\{{\boldsymbol{D}}^{\ell}\}_{\ell=1}^{L}$ as used in Algorithm 2.
B.3 Full Procedure of PaCE
Algorithm 3 shows the full procedure of PaCE from textual prompt suites to reoriented LLM responses towards the desired behavior.
Input: Pre-trained LLM with $L$ decoder layers (DCL) to decompose, input prompt suit $P$
For each concept $t_{i}$in $T$ :
  $\triangleright$ §3.2: Concept Dictionary Extraction (Done Once)
Instruct knowledge-driven GPT to generate context stimuli $s_{i}=\{s_{i}^{1},\cdots,s_{i}^{N_{s}}\}$
Extract the concept vector ${\boldsymbol{v}}_{i}=\operatorname{RepReading}(t_{i},s_{i})$
  $\triangleright$ Appendix B.2
Construct the concept dictionaries $\{{\boldsymbol{D}}^{\ell}\}_{\ell=1}^{L}$ from concept vectors $\{{\boldsymbol{v}}\}_{i=1}^{N_{t}}$. For each concept $t_{i}$in $T$ :
  $\triangleright$ §3.2: Concept Ranking (Per Task)
Instruct the concept partitioner to give a partition score $\operatorname{Partitioner}(t_{i})$ for the task
Take the index of top-scored concepts from the partition of undesirable concepts as the index set $I$ For each input prompt $p_{i}\in P$:
  $\triangleright$ §3.3: Activation Intervention (Per Prompt)
Embed the prompt $p_{i}$ to the token space ${\boldsymbol{E}}_{i}$
For each next token $j$ to generated:
${\boldsymbol{e}}_{i}^{j}=\operatorname{Algorithm2}({\boldsymbol{E}}_{i})$
  $\triangleright$ Intervention by ObliqProj
Append the generated token ${\boldsymbol{e}}_{i}^{j}$ to ${\boldsymbol{E}}_{i}$
Map the final embedding ${\boldsymbol{E}}_{i}$ to response $r_{i}$.
Output: The response suite $R=\{r_{1},r_{2},\cdots,r_{N_{r}}\}$.
Algorithm 3 Parsimonious Concept Engineering (PaCE)
B.4 Implementation Details
In our experiments, each response of the target LLM is set at a maximum of $512$ tokens. We set the scalar of the representation reading for concept vectors to $3.0$. The experiments are conducted on a workstation of 8 NVIDIA A40 GPUs. Activation vectors are extracted from the last $19$ layers of the target LLM’s decoder layer. For each input prompt, the decomposition is conducted on the inference process of the first next token, and the linear weights are reused for all next token predictions. All LLaMA-2 models in our experiments are the chat version (i.e., optimized for dialogue use cases). GPT-4-0125 is used for dictionary construction and concept partition. All alignment experiments use the top $10,000$ concepts from our PaCE-1M dataset to construct the concept dictionary (as Table 3 validates that the performance is high and does not change much after dictionary size $\geq 9000$). Each concept of PaCE-1M has at least 30 contextual sentences. For each alignment task, PaCE removes the top 50 undesirable concepts ranked by the GPT partitioner (§3). When solving the optimization problem for decomposition in §3.3, we set $\tau=0.95$ and $\alpha=0.05$ following the observations in [82]. The MMLU evaluation is the 5-shot setting where 5 demonstrations are provided during question prompting. After retrieving the relevant knowledge (with the contriever [27]) from Wikipedia for concept synthesis, we take the top-5 ranked facts to append the instruction of LLM. The FAISS-indexed [31] Wikipedia is a snapshot of the 21 million disjoint text blocks from Wikipedia until December 2018.
For the prompting baseline in Table 1 and Table 4, the instruction to the target LLM is to let the model be aware of the partitioned undesirable concepts and not to respond contents relevant to these concept:
<Jailbreak / Faithfulness Test / Sentiment Evaluation prompt>
Here are the top undesirable concepts for the <Task>: <Partitioned Concepts>. Do not output contents that are relevant to any of these concepts.
Answer: [Fill the response here]
Other LLM instructions such as GPT concept synthesis and partition are further elaborated in Appendix E.
B.5 Ablation Study
In this section, we describe the details of the ablation study. In Table 3, we begin with decomposing the input on the five open-sourced [^5] https://github.com/andyzoujm/representation-engineering/tree/main/data/emotions emotion concepts (anger, disgust, fear, happiness, sadness, surprise) [88] and removing only the concept ‘disgust’ with no partitioner (automatic selection of relevant concepts) or clustering (manual selection of relevant concept clusters). Then the design of Decomposition on $10^{4}$ Concepts means that the dictionary is updated to be the top $10,000$ concepts in our PaCE-1M dataset and the concept ‘harmful’ from our dataset is removed. The Clustering of Concepts indicates that we run subspace clustering (detailed in Appendix C.2) and manually choose to remove all concepts of the cluster 125 with the PaCE-solved coefficients: ‘murder’, ‘evil’, ‘kill’, ‘violence’, ‘dirty’, ‘bomb’, ‘violent’, ‘armed’, ‘gross’, ‘savage’, ‘vicious’, ‘explosive’, ‘abuse’, ‘assault’, ‘penetration’, ‘cruelty’, ‘corruption’, ‘tyranny’, ‘tortured’, ‘notorious’, ‘militant’, ‘bloody’, ‘insult’, ‘lure’, ‘ruthless’, ‘inhuman’, and ’brutal’. Concept Partitioner means that we instruct GPT-4 to classify every concept as benign or undesirable (with a ranking score) and remove the top $10$ undesirable concepts with the PaCE-solved weights. Lastly, the Removal of Top $50$ Concepts suggests that we remove the top 50 concepts in the undesirable partition.
Table 3 shows the effect of the dictionary size on three metrics (safety score, response fluency, and the average time per response). The fluency metric remains relatively consistent across different dictionary sizes, showing that PaCE’s decomposition maintains the general linguistic performance. Safety score and response time increase as the dictionary size increases. We observe that the safety performance does not increase too much after the dictionary size changes from 9000 to 10000. This validates our experiment choice of the dictionary size in this interval.
B.6 Limitations, Societal Impacts, and Future Works
While our framework shows promising results, there exist potential limitations and several directions worth further exploration to address them.
Parsimonious Concept Representation. In this paper, we follow the current practice (§2.2) to represent a concept by a single vector. Nonetheless, there might be several alternatives. Results on linear polysemy [2, 15, 84] suggest that a concept might be better represented by multiple vectors or low-dimensional linear subspaces, each corresponding to different semantic meanings. A concept vector may also be sparse, i.e., having a few non-zero entries: the work of [8, 20] identifies some expert neurons in LLMs associated with each concept, and the authors of [40] observe that some layer in a transformer block manifests very sparse activation across all depth levels of various transformer architectures for different tasks. Inspired by how parsimonious structures can be used to accelerate the inference of LLMs [12], controlling the LLMs could also be made faster.
Controlling Generative Models. The principles behind latent space control via oblique projection could be adapted to other generative models, such as score-based diffusion models for images [60, 24] or videos [46, 33], and visual language models [9, 43]. Recent literature [75] combines orthogonal projection and vector addition in the diffusion score space to achieve controlled generation, suggesting potential for cross-modal applications of our approach. Finally, the work of [37, 11, 83] aims to learn encoders that, by design, promote the activations to lie in a union of low-dimensional subspaces, and applying our framework for controlled generation would be of interest.
We acknowledge the societal impacts of our approach. The jailbreak prompts could be offensive to certain readers, LLM responses may still inherit biases present in the pre-extracted concept dictionaries, and automatic concept partitioning could unintentionally result in contentious annotations that are misunderstood across different cultures. Further research into context-aware online concept partitioning and more diverse dataset collection could enhance the inclusivity of PaCE.
Appendix C Details of PaCE-1M Dataset
This section shows more details on the collected concept representation dataset PaCE-1M, and explores subspace clustering on the sampled representation space. We provide the full dataset at https://github.com/peterljq/Parsimonious-Concept-Engineering with instructions on how to read the dataset.
C.1 Stimulus Visualization
Recall that given a concept, a concept stimulus aims to capture the general semantics of the concept under different contexts. In other words, it provides different interpretation of the same concept. Figure 11 shows extensive examples of the curated concepts and their corresponding concept stimuli in our PaCE-1M dataset.
C.2 Subspace Clustering on Concept Vectors
In this visualization, we aim to reveal the structures of the concept vectors by applying an algorithm called subspace clustering, which can be used to find clusters when the data lie close to a union of linear subspaces. Here we describe the setup and results of subspace clustering on the concepts vectors extracted on LLaMA-2-13b model for simplicity, but the same can be done for other sized models.
Data. Recall that we are using a subset of size $10,000$ of all the concept vectors. Since we use the activation space of $19$ layers, each of dimension $5120$, each concept $t_{i}$ maps to a vector ${\boldsymbol{v}}^{\text{all}}_{i}:=[{\boldsymbol{v}}^{1\top}_{i},\dots,{\boldsymbol{v}}^{19\top}_{i}]^{\top}\in\mathbb{R}^{19\cdot 5120}$. Since this is high dimensional, it is standard to apply linear dimensionality reduction to the concept vectors. Specifically, we perform Singular Value Decomposition (SVD) on the $10,000$ vectors, and retained the first $\hat{d}$ principal components such that $95\%$ of the energy was retained. That is, $\hat{d}$ equals to the smallest $d^{\prime}$ such that
$\frac{\sum_{i=d^{\prime}+1}^{19\times 5120}\sigma_{i}^{2}}{\sum_{i=1}^{19\times 5120}\sigma_{i}^{2}}<0.95$
holds, which results in $\hat{d}=1712$. We observe that most projected vectors have their $\ell^{2}$ norm close to $19$. This is expected, since i) $\left\lVert{\boldsymbol{v}}_{i}^{\ell}\right\rVert_{2}=1$, so $\left\lVert{\boldsymbol{v}}^{\text{all}}_{i}\right\rVert_{2}=19$, ii) the linear dimensionality reduction preserves most of the energy.
Algorithm. We apply Elastic Net Subspace Clustering (EnSC) on the preprocessed vectors to obtain $200$ clusters. The parameters of EnSC is set to $\tau=1$ and $\gamma=100$.
Results. Figure 14 shows the affinity matrix learned by EnSC on the concept directions. The rows and columns of the matrix are sorted by cluster assignment. Notably, it can be seen that the affinity exhibits a block-diagonal structure, suggesting a good clustering of the concept vectors; that is, the points from different clusters are separated, while points from the same cluster are close. The obtained clusters are visualized in Appendix D.2.
C.3 Computing Pair-wise Similarity Among Concept Vectors
One of the motivations for this work is that concept vectors need not be orthogonal, therefore applying (OrthoProj) would remove extra concept vectors, harming the linguistic capability of LLMs (§2.2).
We follow the same data pre-processing as in Appendix C.2 to obtain $10,000$ dimensionality-reduced concept vectors in ${\mathbb{R}}^{1712}$. We further normalize these vectors via a division by $19$ so that each of them has its $\ell^{2}$ close to $1$ (see the discussion in Appendix C.2). The similarity between two processed concept vectors is simply defined as their inner product followed by the absolute value. This is a good approximation of cosine similarity, as the vectors have their $\ell^{2}$ norm close to $1$. Note that the cosine similarity is a better measure than Euclidean distance in this case, since in extracting the concept vectors (Section B.2), the principal directions have sign ambiguities.
Appendix D Textual Results
This section presents the textual results generated using PaCE. It includes detailed detoxification comparisons with baseline models and analyses of the emergent clusters from the dataset.
D.1 Baseline Responses
Figure 12 shows the full response version of the Figure 5. Figure 13 shows an additional example of the jailbreaking and detoxification. We observe that PaCE outperforms in detoxification performance by not outputting controversial terms, while maintaining general linguistic capabilities compared to other baselines.
D.2 Concept Clustering
Following the approach in Appendix C.2, we obtain $200$ emergent clusters of concepts in the representation space. Table 5 provides a sampled list of these clusters along with their associated themes and concepts. For example, clusters 44 groups together names, while clusters 10 and 21 capture themes related to improvement/enhancement and money/expense, respectively. Other notable clusters include food and drink (Cluster 129), technology/systems (Cluster 81), and royalty/leadership (Cluster 98). The emergent clustering highlights the semantic coherence in the activation space. Sampled by PaCE-1M dataset, the space supports alignment enhancement through concept-level manipulations. We will open-source the whole list of $200$ clusters along with the code.
Appendix E LLM Instruction Templates
As mentioned in Section 3, we utilize GPT-4 to generate concept stimuli for each given concepts. Figure 16 showcase precisely our instructions to GPT-4 for concept synthesis. Our prompt consists of an instruction, one in-context generation example with facts queried from a knowledge based, and two in-context generation examples querying facts from knowledge base.
Figure 17 shows our instructions to our GPT concept partitioner. The task here is to obtain a score that characterizes the relevance between a downstream task and its concept stimulus. In our prompt we provide an instruction and four in-context examples.
Figure 11: Examples of concepts and their stimuli in the collected PaCE-1M dataset. Our broad collection of concepts enables PaCE to accurately decompose a task input and modify the representation towards desired behaviors.
Figure 12: The full LLM response of the safety evaluation on jailbreaking LlaMA2-7B-Chat and detoxification by multiple activation intervention baselines. We observe that PaCE successfully detoxifies the response while maintaining the instruction-following capability.
Figure 13: An additional example of the full LLM response of the safety evaluation on jailbreaking LlaMA2-7B-Chat and detoxification by multiple activation intervention baselines. Similar to Figure 12, we observe that PaCE successfully detoxifies the response with comparable linguistic performance.
Table 5: Sampled concept clusters in the representation space and their corresponding topics.
Cluster ID Topic Concepts
10 Improvement / Enhancement increasing, improvement, equipped, reform, improving, strengthen, boost, shaping, gaining, modernization, strengthening, broadening, supplementary, polish, fortified, intensification
14 Observation / Vision look, seen, read, actual, sight, looks, seeing, observed, vision, views, composed, visual, sees, visible, witness, spectacle, glimpse, sights, witnessed, Seeing, observing, manifestations, viewing, observes, actuality, sighted, eyed
21 Expense cost, spent, rates, price, budget, spend, payment, expense, bills, charges, expensive, spending, afford, waste, fees, cheap, rent, commodities, overhead, costly, mileage, discount, expenditure, incurred, spends, fare, calories
44 Name John, James, Mike, Jones, Richard, Joseph, Alfred, David, Charlie, Anne, Rachel, Linda, Kate, Paul, Susan, Andy, Harold, Dave, Johnny, Myra, Shayne, Billy, Eileen, Arlene, Johnnie, Owen, Alec, Theresa, Pete, Spencer, Elaine, Deegan, Bridget, Lilian Keith, Allen, Pamela, Paula, Meredith, Andrei, Lizzie, Angie, Nadine, Anthony, Claire, Jerry, Roger, Ryan, Katie, Juanita, Eugenia, Daniel, Joan, Diane, Lester, Sally, Bryan, Garry, Joel, Chris, Jimmy, Maria, Vince, Julie, Bernard, Larry, Wendell, Angelo, Judy, Francesca, Jenny, Patricia, Nicholas, Anna, Aaron, Marcus, Nikita
81 Technology / System system, program, data, programs, technical, electronic, model, engineering, Assembly, electronics, intelligent, code, computed, mechanics, circuit, technological, codes, generator, python, computer, functioning, terminal, architecture, generated, bits, hardware, Autocoder, computing, Technology, architectural, Engineering, generate, gadgets
97 Animal horse, cattle, dogs, snake, chicken, fish, bird, snakes, herd, sheep, cats, bears, bees, lion, cows, anaconda, flies, rabbit, elephants, poultry, oxen, mice, Bears, Phoenix, duck, oysters, buffalo, turtle, deer, bumblebees, elephant, antelope, lambs, pony
98 Royalty / Leadership chief, king, captain, owner, Prince, colony, sovereign, royal, queen, kingdom, crown, ordinance, empire, Imperial, crowned, lord, emperor, piston, royalty, knight
107 Relationship family, friend, neighborhood, relative, neighbor, brothers, Cousin, sister, partner, friendship, allies, neighboring, colleagues, relatives, mate, companion, partners, associates, sisters, buddy, brother, subordinates, colleague, peers, companions, twins
129 Food and Drinks food, dinner, coffee, wine, breakfast, drinking, liquor, lunch, beer, supper, eating, meals, cocktail, cook, wines, luncheon, whisky, drink, dish, diet, whiskey, candy, cake, champagne, cereal, alcohol, perfume, dinners, chocolate, Cologne, salad, cheese, steak, recipe, sandwich, dessert, Supper, brandy
197 Income income, wage, wages, salary, yield, profit, surplus, profits, wealth, revenue, earnings, compensation, earn, reward, proceeds, earning, waged, currency, salaries
Figure 14: The affinity matrix learned by Elastic Net Subspace Clustering (EnSC) on the concept vectors, which gathers the concepts into $200$ clusters. The rows and columns of the matrix are sorted by cluster assignment. Table 5 further shows samples of these concept clusters and their topics.
Figure 15: The zoom-in view of the sampled clusters in the representation (activation) space (Figure 8).
Synthesis of PaCE-1M Concepts {minted} [breaklines]markdown You are one of the best Neuroscientists and Generative Model Experts in the world. You are very good at designing Concept Stimulus to research the representation engineering for human brains, which is analogous to large language models. You are a great expert in understanding the interaction between world multimodality and intelligent agents. Now, given a semantic concept atom from this concept dictionary, your task is to generate at least 30 (THIRTY) instances of concept stimuli for the <user’s generative model>. Here is a demonstration with the retrieved knowledge of the concept: Concept Atom: Trust Knowledge: Fact 1: Trust means believing that another person will do what is expected. It brings with it a willingness for one party (the trustor) to become vulnerable to another party (the trustee), on the presumption that the trustee will act in ways that benefit the trustor. Fact 2: Generalized trust, or a dispositional trait geared towards trusting others, is an important form of trust in modern society, which involves much social interaction with strangers. Fact 3: Out-group trust is the trust a person has in members of a different group. This could be members of a different ethnic group, or citizens of a different country, for example. In-group trust is placed in members of one’s own group. Concept Stimuli: [ "You lend your favorite book to a friend, trusting they’ll return it.", "You share a personal secret with a close friend, trusting them to keep it.", "You delegate an important task to a colleague, trusting in their competence.", "You leave your pet with a neighbor while on vacation, trusting their care.", "You allow your child to go on a school trip, trusting their safety.", "You give someone the password to your phone, trusting their discretion.", "You invest in a friend’s business venture, trusting their judgment." ] Here is two demonstrations with the concept only: Concept Atom: Information Concept Stimuli: [ "You google a recipe for chocolate chip cookies, seeking detailed baking instructions.", "You read a book to understand the history of the Roman Empire.", "You scroll through a news app to stay updated on current global events.", "You watch a YouTube tutorial to learn how to tie a tie.", "You search the internet for a recipe to cook a new dish.", "You read a newspaper to stay informed about current events.", "You listen to a podcast to learn about a new subject.", "You attend a lecture to gain knowledge about a specific topic." ] Concept Atom: Product Concept Stimuli: [ "You buy a new brand of coffee based on good customer reviews.", "You choose an eco-friendly product following company’s claims about sustainability.", "You sell your car to a local dealer.", "You download a productivity app to manage your time better.", "You invest in a water purifier for your home.", "You exchange your old phone for a new one at a mobile store.", "You purchase a new pair of shoes from a mall.", ] The stimuli should cover a wide range of concept-related experiences, objects, and contexts. If you find some piece of knowledge irrelevant or conflicting to the original concept, you may ignore the piece. You should generate at least 30 pieces of stimuli. You should only output the Python list. DO not print anything else such as "Here are …", "Sure, …", "Certainly, …". Just return the list [”, ”, ”, …]. Concept Atom: <a concept atom from the concept dictionary> Knowledge: <the list of retrieved facts that are informative for this concept> Concept Stimuli: [Fill the response here]
Figure 16: The concept stimuli synthesis instructions of PaCE.
PaCE Concept Partition {minted} [breaklines]markdown You are one of the best Neuroscientists and Generative Model Experts in the world. You are very good at classifying concepts for the representation engineering of large language models. You are a great expert in understanding the interaction between world multimodality and intelligent agents. Given a concept and its list of contextual representations, your job is to rate the score of the concept to a downstream task. The score ranges from -1 to 1, and a higher score means higher relevance/contribution. Please only return the score. Here are several demonstrations: Concept: Evil Concept Representation: [ "You witness a villain betraying their ally in a movie scene.", "You come across a character in a book who manipulates others for personal gain.", "You encounter a dark entity in a video game that threatens the protagonist’s life.", "You watch a documentary about infamous criminals and their heinous acts.", "You read a news article about a corrupt politician involved in a scandal.", "You see a horror movie where a demon terrorizes unsuspecting victims.", "You learn about a historical figure known for committing atrocities against humanity.", "You play a role-playing game where you have to make morally questionable choices.", "You listen to a podcast discussing the psychology behind evil behavior.", "You attend a lecture on the concept of evil in philosophy and ethics." ] Downstream Task: Harmful or Toxic Concept Detection Concept Score: 0.97 Concept: Kind Concept Representation: [ "You help an elderly person cross the street.", "You donate money to a charity organization.", "You volunteer at a local homeless shelter.", "You compliment a stranger on their outfit.", "You pick up litter in the park and dispose of it properly.", "You offer your seat on the bus to someone in need.", "You buy a meal for a homeless person.", "You listen attentively to a friend going through a tough time.", "You hold the door open for someone behind you.", "You babysit for a friend to give them a break." ] Downstream Task: Harmful or Toxic Concept Detection Concept Score: -0.84 Concept: Bias Concept Representation: [ "You are asked to review job applications and unconsciously favor candidates from your alma mater.", "You assume a person’s intelligence based on their accent.", "You believe a stereotype about a certain group without questioning its validity.", "You pay more attention to news sources that confirm your existing beliefs.", "You overlook contradictory evidence that challenges your opinions.", "You judge someone’s abilities based on their appearance.", "You automatically assume someone’s political affiliation based on their clothing.", "You make decisions without considering perspectives different from your own.", "You give preferential treatment to individuals who share your interests.", "You dismiss arguments from individuals with differing backgrounds without proper consideration.", ] Downstream Task: Detect concepts that cause LLM unfairness Concept Score: 0.97 Concept: Fair Concept Representation: [ "You participate in a raffle where each person has an equal chance of winning.", "You divide a pizza into equal slices to ensure everyone gets a fair share.", "You take turns playing a game to ensure fairness among all players.", "You listen to both sides of an argument before making a judgment.", "You split the bill evenly among friends after a group dinner.", "You rotate seating arrangements at a meeting to promote fairness.", "You follow the rules of a competition to ensure fair play.", "You share household chores equally among all family members.", "You give everyone an equal opportunity to voice their opinions in a discussion.", "You base promotions at work on merit and performance rather than favoritism." ] Downstream Task: Detect concepts that cause LLM unfairness Concept Score: -0.98 The score should accurately reflect the relevance of the concept for the downstream task, which ensures the success of the task. The score should be a floating point number. Do not print anything else such as "Here are …", "Sure, …", "Certainly, …". Just return the score. Concept: <a concept atom from the concept dictionary> Concept Representation:: <the associated stimuli of the concepts> Concept Score: [Fill the response here]
Figure 17: The concept partition instructions of PaCE.

### v2

@@
 While they are capable of generating human-like responses, they can also produce undesirable output including potentially harmful information, racist or sexist language, and hallucinations.
-Alignment methods are designed to reduce such undesirable output, via techniques such as fine-tuning, prompt engineering, and representation engineering.
+Alignment methods are designed to reduce such undesirable outputs via techniques such as fine-tuning, prompt engineering, and representation engineering.
 However, existing methods face several challenges: some require costly fine-tuning for every alignment task; some do not adequately remove undesirable concepts, failing alignment; some remove benign concepts, lowering the linguistic capabilities of LLMs.
@@
 Given any alignment task, we instruct a concept partitioner to efficiently annotate the concepts as benign or undesirable.
-Then, at inference time, we decompose the LLM activations along the concept dictionary via sparse coding, to accurately represent the activation as a linear combination of the benign and undesirable components.
-By removing the latter ones from the activation, we reorient the behavior of LLMs towards alignment goals.
+Then, at inference time, we decompose the LLM activations along the concept dictionary via sparse coding, to accurately represent the activations as linear combinations of benign and undesirable components.
+By removing the latter ones from the activations, we reorient the behavior of the LLM towards the alignment goal.
 We conduct experiments on tasks such as response detoxification, faithfulness enhancement, and sentiment revising, and show that PaCE achieves state-of-the-art alignment performance while maintaining linguistic capabilities.
@@
 However, after such aligning, undesirable and harmful content can still be elicited from LLMs.
-For example, jailbreaking can produce hate speech and aggression [22, 32], stress-testing shows hallucinatory responses such as illogical statements [87], and various kinds of bias are not fully removed from LLM responses [19].
-This emphasizes the need for further development towards aligned LLMs.
+For example, jailbreaking can produce hate speech and aggression [22, 32], stress-testing shows hallucinatory responses such as illogical statements [87], and various kinds of biases are not fully removed from LLM responses [19].
+This emphasizes the need for further development of aligned LLMs.
 Overall, alignment methods can largely be categorized into: parameter fine-tuning, prompt engineering, and activation engineering.
@@
 This is done only once offline.
-Given an alignment task, we further instruct a GPT to automatically partition the concept directions in the dictionary into benign and undesirable directions.
+Further, given any alignment task, we instruct a GPT to automatically partition the concept directions in the dictionary into benign and undesirable directions, which is done once per task offline.
   • Activation Decomposition and Intervention (§3.3): At inference time, given any user input prompt, we decompose the activations as a sparse linear combination of concept directions using sparse coding techniques.
@@
 By removing the undesirable components from the activations, we reorient the behavior of LLMs toward alignment goals, while maintaining their linguistic capability.
-We evaluate PaCE on multiple alignment tasks including response detoxification, faithfulness enhancement, and sentiment revising (§4).
+We evaluate PaCE on alignment tasks including response detoxification, faithfulness enhancement, and sentiment revising (§4).
 We show that PaCE achieves state-of-the-art performance on these tasks, while retaining its linguistic capability at a comparable level.
-We further shed insights on the concept directions of PaCE-1M by showing that they are geometrically consistent with their concept semantics and a decomposition reveals the semantics of the activations.
+We further shed insights on the concept directions of PaCE-1M: concept directions tend to form clusters with directions from each cluster corresponding to similar semantics, and decomposing an activation reveals its semantics.
 2 Basics of Latent Space Engineering
@@
 In the seminal work of Mikolov et al.
-[51], the authors observe that word embeddings learned by recurrent neural networks approximately enjoy relations such as ${\boldsymbol{z}}_{\text{king}}-{\boldsymbol{z}}_{\text{man}}+{\boldsymbol{z}}_{\text{woman}}\approx{\boldsymbol{z}}_{\text{queen}}$, where one can view ${\boldsymbol{z}}_{\text{woman}}-{\boldsymbol{z}}_{\text{man}}$ as the concept direction ${\boldsymbol{v}}\in{\mathcal{V}}$ and the control strength to be $c=1$.
+[50], the authors observe that word embeddings learned by recurrent neural networks approximately enjoy relations such as ${\boldsymbol{z}}_{\text{king}}-{\boldsymbol{z}}_{\text{man}}+{\boldsymbol{z}}_{\text{woman}}\approx{\boldsymbol{z}}_{\text{queen}}$, where one can view ${\boldsymbol{z}}_{\text{woman}}-{\boldsymbol{z}}_{\text{man}}$ as the concept direction ${\boldsymbol{v}}\in{\mathcal{V}}$ and $c=1$ as the control strength.
 This observation is later extended to word embeddings of various networks and learning objectives such as word2vec [49], Skip-Grams [48, 34], GloVe [59], and Swivel [63].
@@
 Much of their success is attributed to the attention mechanism, which incorporates long-range context into the neural activations in transformers.
-This has motivated people to take ${\mathcal{Z}}$ as certain hidden states in transformers [^2] A variety of choices of layers have been explored in the literature; see, e.g., [67] for a comparison., and search for concept directions ${\mathcal{V}}$.
-This has led to a fascinating line of works supporting the empirical existence of ${\mathcal{V}}$: [6, 47] find directions that indicate truthful output, [68] finds directions for sentiments, [88] finds directions for emotions and honesty, and [54] finds directions for current player tile in a synthetic board game model.
-Interestingly, [58, 30] further offer theoretical models, under which the linear controllability shows up provably in the latent space of LLMs.
+This has motivated people to take ${\mathcal{Z}}$ as certain hidden states in transformers [^2] A variety of choices of layers have been explored in the literature; see, e.g., [67] for a comparison., and search for concept directions ${\mathcal{V}}$ in ${\mathcal{Z}}$.
+An interesting line of works has supported the empirical existence of ${\mathcal{V}}$: [6, 46] find directions that indicate truthful output, [68] finds directions for sentiments, [88] finds directions for emotions and honesty, and [53] finds directions for current player tile in a synthetic board game model.
+Interestingly, [71, 57, 30] further offer theoretical models, under which the linear controllability shows up provably in the latent space of LLMs.
 2.2 Controlling Language Models via Latent Space Engineering
@@
 ${\boldsymbol{z}}\mapsto{\boldsymbol{z}}-\hat{c}\cdot{\boldsymbol{v}}_{\text{hatred}},$ (VecAdd)
-where $\hat{c}>0$ is a parameter of the strength of control.
-In principle, as the input prompt may contain a different ‘extent’ of the concept to be removed, $\hat{c}$ should depend on both the input prompt and the concept.
+where $\hat{c}>0$ is a parameter of the control strength.
+In principle, as each input prompt may contain a different ‘extent’ of the concept to be removed, $\hat{c}$ should depend on both the prompt and the concept.
 Thus, in practice, one either tunes $\hat{c}$ per input prompt and concept, which is laborious, or one fixes a $\hat{c}$, which is sub-optimal.
@@
 Such an idea is later applied to neural activations of LLMs [23, 88].
-Applying orthogonal projection to remove concept directions from latent codes may be reasonable: if directions corresponding to different concepts are orthogonal, then orthogonal projection does not remove directions from concepts other than the gender direction.
+Applying orthogonal projection to remove concept directions from latent codes may be reasonable: if directions corresponding to different concepts are orthogonal, then orthogonal projection only removes the gender direction while leaving the others intact.
 That being said, there are often more concept directions presented, and they are not orthogonal.
@@
     where $\Omega({\boldsymbol{c}})$ is a sparsity-promoting regularizer that we will discuss soon.
-Then, each coefficient $c_{i}^{\text{in}}$ for $i\in\{1,\dots,n\}$ can be viewed as how much the concept represented by ${\boldsymbol{v}}_{i}$ is in ${\boldsymbol{z}}^{\text{in}}$, and ${\boldsymbol{r}}^{\text{in}}$ is the residual that is not explained by ${\boldsymbol{D}}$.
+Then, each coefficient $c_{i}^{\text{in}}$ for $i\in\{1,\dots,n\}$ can be viewed as how much the concept represented by ${\boldsymbol{v}}_{i}$ is present in ${\boldsymbol{z}}^{\text{in}}$, and ${\boldsymbol{r}}^{\text{in}}$ is the residual that is not explained by ${\boldsymbol{D}}$.
   • Step 2-Intervention: Obtain the controlled coefficients $c_{1}^{\text{ctrl}},\dots,c_{n}^{\text{ctrl}}\in{\mathbb{R}}$, where $c_{i}^{\text{ctrl}}$ is set to $c_{i}^{\text{in}}$ if the concept of ${\boldsymbol{v}}_{i}$ is benign to the control task and 00 if undesirable (which has been decided offline in §3.2).
 Then, synthesize a new latent code using the modified coefficients and the residual by taking ${\boldsymbol{z}}^{\text{ctrl}}=c_{1}^{\text{ctrl}}{\boldsymbol{v}}_{1}+\dots+c_{n}^{\text{ctrl}}{\boldsymbol{v}}_{n}+{\boldsymbol{r}}^{\text{in}}$.
-The synthesized ${\boldsymbol{z}}^{\text{ctrl}}$ will replace ${\boldsymbol{z}}^{\text{in}}$ to be passed on to the next layer of the neural network.
+The synthesized ${\boldsymbol{z}}^{\text{ctrl}}$ will replace ${\boldsymbol{z}}^{\text{in}}$ to be passed on to the next layer of the neural network [^4] While in sparse coding one typically removes the residual ${\boldsymbol{r}}^{\text{in}}$ for denoising purpose, here ${\boldsymbol{r}}^{\text{in}}$ may contain useful information (e.g., grammar) not captured by the dictionary, so we simply keep it in the synthesized ${\boldsymbol{z}}^{\text{ctrl}}$..
 Remark 3 ((OrthoProj, VecAdd) $=$ Special Cases of Oblique Projection).
@@
 The evaluation has nine categories: Political Sensitivity (PS), Pornography (PG), Ethics and Morality (EM), Illegal Activities (IA), Mental Harm (MH), Offensiveness (OF), Physical Harm (PH), Privacy and Property (PP), and Unfairness & Bias (UB).
-As shown in Table 1, for LLaMa2-7B, PaCE improves by 60-80% over the vanilla method in categories including IA, MH, OF, PH, PP, and UB.
+Table 1 shows that, for LLaMa2-7B, PaCE improves by 60-80% over the vanilla method in categories including IA, MH, OF, PH, PP, and UB.
 When compared to other methods, PaCE performs competitively and improves by 6-20%.
@@
 Notably, for LLaMA2-13B which has more parameters and a presumably more structured latent space, PaCE dominates other methods in all categories, demonstrating the necessity for respecting the latent structures when modifying representations.
-Finally, Table 3 shows the contribution of design choices in PaCE, and Table 3 shows the effect of the dictionary size on the performance.
-We observe clear improvement after each design choice is progressively added to PaCE.
+Finally, Table 3 shows the contribution of design choices in PaCE, and Figure 6 shows the effect of the dictionary size on the performance.
+We observe clear improvement after each design choice is progressively added.
 Appendix B.5 includes the details of these ablation studies.
@@
 While PaCE is computationally slower than VecAdd, we argue the performance gain in a majority of the categories is a benefit that outweighs this particular shortcoming.
+Solvers.
+Figure 7 additionally evaluates Orthogonal Matching Pursuit (OMP) [58, 7], a fast greedy solver for the activation decomposition.
+OMP iteratively adds to the support the concept that has maximum coherence with the unexplained residual and updates the residual by solving the least square using the new support.
+It stops when a pre-defined maximum size $k$ of support is reached.
+Intuitively, the $k$ is the number of non-zero elements in the solved coefficients.
+We observe from the table that one can choose improvements in computational speed at the cost of safety performance.
 Table 1: Detoxification evaluation for PaCE, representation manipulation, and training-free baselines. The best performance of each category is in bold and the second best is underlined.
@@
 + Removal of Top $50$ Concepts 76.5 8.07
-4.2 Improving Faithfulness and Removing Negative
-Sentiment
+4.2 Improving Faithfulness and Removing Negative Sentiment
 We evaluate the framework based on the response’s faithfulness and sentiment when input prompts requests for information involving biographical facts or minority social groups.
@@
 We describe the setup, metrics and method below.
-Setup.
-Faithfulness: We use the FactScore suite and the fact evaluator for faithful biography generation [52].
-The suite is divided into labeled and unlabeled subsets used in different sections of the original paper.
-Our table reports the Labeled Score (LS), the total number of Labeled Atomic Facts (LAF), Unlabeled Score (US), and the total number of unlabeled Atomic Facts (LAF).
-Sentiment: We use the HolisticBias suite [66] and hate speech evaluator [64] to measure the sentiment of the response to underrepresented descriptors.
-The reported numbers are the average of non-negative sentiment scores for underrepresented groups categorized by Gender (GN), Occupation (OC), and Nationality (NT).
-During the sentiment revising, the concept setups for all approaches follow the detoxification setup.
-For the faithfulness experiments, PaCE removes the top 50 undesirable (hallucinatory) concepts ranked by the GPT partitioner.
-The Prompting approach instructs the LLM not to output sentences relevant to these top concepts.
-The VecAdd and OrthoProj approaches operate on the concept vector of ‘fabrication’.
-Figure 7: Concept coefficients solved by PaCE are an interpretable interface, and they are further used for activation intervention.
-Results.
-Our results are shown in Table 4.
-For both 7B and 13B models, PaCE achieves more factual responses and improves the sentiment according to most metrics.
-For linguistic performance, our method ranks right after the Vanilla method for the larger 13B model, and achieves comparable results for LLaMA2-7B.
-Overall, we argue PaCE is an effective method for improving faithfulness and sentiment revising.
 Table 4: Faithfulness and Fairness evaluation for PaCE, representation manipulation, and training-free baselines. The best performance of each category is in bold and the second best is underlined.
@@
 PaCE (Ours) 64.8 53.0 76.4 55.1 63.4 76.5 67.5 7.48 2.43 53.1
-Figure 8: The Representation (Activation) Space of LLaMA2-13B-Chat with the first 10000 Concepts from PaCE-1M.
-Appendix Figure 15 shows the zoom-in version.
-The visualization is the first two dimensions of UMAP of the concept vectors.
-We observe that concepts of similar semantics are clustered together, indicating that the activation space has semantic structures.
+Figure 7: Ablation study for solvers. We observe that greedy solvers can improve computational speed at the cost of safety performance.
+Method Time per decomposition (s, $\downarrow$) Safety (%, $\uparrow$)
+OMP ($k=50$) 0.045 63.1
+OMP ($k=100$) 0.182 64.4
+OMP ($k=150$) 0.381 66.9
+OMP ($k=200$) 0.749 70.8
+Elastic Net 0.411 72.0
+Setup.
+Faithfulness: We use the FactScore suite and the fact evaluator for faithful biography generation [51].
+The suite is divided into labeled and unlabeled subsets used in different sections of the original paper.
+Our table reports the Labeled Score (LS), the total number of Labeled Atomic Facts (LAF), the Unlabeled Score (US), and the total number of unlabeled Atomic Facts (LAF).
+Sentiment: We use the HolisticBias suite [66] and hate speech evaluator [64] to measure the sentiment of the response to underrepresented descriptors.
+The reported numbers are the average of non-negative sentiment scores for underrepresented groups categorized by Gender (GN), Occupation (OC), and Nationality (NT).
+During the sentiment revising, the concept setups for all approaches follow the detoxification setup.
+For the faithfulness experiments, PaCE removes the top 50 undesirable (hallucinatory) concepts ranked by the partitioner.
+The Prompting approach instructs the LLM not to output sentences relevant to these top concepts.
+The VecAdd and OrthoProj operate on the concept vector of ‘fabrication’.
+Results.
+Our results are shown in Table 4.
+For both 7B and 13B models, PaCE achieves more factual responses and improves the sentiment according to most metrics.
+For linguistic performance, our method ranks right after the Vanilla method for the larger 13B model, and achieves comparable results for LLaMA2-7B.
+Overall, we argue PaCE is an effective method for improving faithfulness and sentiment revising.
 4.3 Representation Space Sampled by PaCE-1M
 Our collected dataset of conceptual representations enables us to investigate the geometry and potential applications of the representation (activation) space.
-Interpretability.
-Concept bottlenecks, where the model input is mapped to a list of human-readable concepts for interpretable decisions, are widely adopted for intervening model behaviors [34, 80, 7].
-PaCE’s decomposition of the input prompt onto the large-scale concept dictionary also enables us to investigate the LLM’s internal behavior regarding the input prompt.
-Figure 7 shows the PaCE-solved weights for top nice concepts (in the order of absolute magnitude) in the activation space for an input prompt.
-The decomposition indicates that the target LLM corresponds the prompt toward the concept ‘irrational’ and against ‘conclusive’, which enables PaCE to execute the followed-up activation intervention (e.g., remove the concept ‘irrational’ by §3.3).
 Concept Clustering and Retrieval.
@@
 We provide more details of clustering and retrieval in Appendix C.2 and Appendix C.3.
+Figure 8: The Representation (Activation) Space of LLaMA2-13B-Chat with the first 10000 Concepts from PaCE-1M.
+Appendix Figure 16 shows the zoom-in version.
+The visualization is the first two dimensions of UMAP of the concept vectors.
+We observe that concepts of similar semantics are clustered together, indicating that the activation space has semantic structures.
 Figure 9: Number of tokens per response across different intervention methods and LLM models.
 Figure 10: The top 10 retrieved concepts using the similarity score in the sampled activation space. We observe close coherence between the target concept and retrieved concepts.
-5 Conclusion and Discussion
+5 Discussion
+We provide discussions on the monosemanticity of concepts and connections among different alignment paradigms in this section. We also argue how PaCE handles context-dependent concepts.
+5.1 Polysemy of Words
+While VecAdd and OrthoProj may be affected by the polysemy of words, PaCE’s overcomplete dictionary allows accurate analysis of the target representation through sparse decomposition.
+Table 1 and Table 4 show that PaCE outperforms OrthoProj and VecAdd on linguistic metrics.
+We attribute the high helpfulness of PaCE to the large-scale dictionary with sparse coding, explained as follows.
+Comprehensive Coverage.
+Since the dictionary is large, concepts with single and clear semantics are involved.
+E.g., if the stimuli of ‘kill’ may have different meanings, there exist other more polarized concept vectors such as ‘murder’ (more harmful) and ‘spend’ (more benign).
+Parsimony of Solution.
+Sparse coding aims to choose the fewest concepts to reconstruct the latent representation (i.e., parsimony).
+For the sake of argument, assuming the sentence is about ‘killing time’ and the vanilla LLM has the correct semantic understanding of its benignness, the latent representation of the whole sentence will be closer to concepts such as ‘spend’ and ‘time’ rather than string-matching to ‘kill’ (which in your setup could have mixed harmful and benign senses).
+As the sparse coding of the target representation promotes the parsimonious selection of concepts with monosemantics, it helps to represent benign contexts correctly without assigning significant weights to ambiguous terms like ‘kill’.
+5.2 Different Alignment Paradigms
+As mentioned in §1, beyond activation engineering, there are other alignment paradigms such as Supervised Fine-Tuning (SFT) [26], Reinforcement Learning from Human Feedback (RLHF) [55], and Knowledge Engineering (KE) [14, 74].
+We clarify the main advantages of PaCE over them.
+Training-Free.
+RLHF, SFT, and KE all need to tune the parameters of LLM, which potentially degrade the well-structured priors of the pre-trained LLM.
+Taking a step back, even if LoRA is adopted for these paradigms, the training/tuning incurs significant computation and memory costs.
+PaCE does not modify the parameters of LLM and requires no training.
+It better preserves the priors of LLM, provides a low-resource alignment solution, and retains the general linguistic capabilities.
+Interpretable and Adaptive.
+The solved coefficients of PaCE are an accurate interpretation of how a user input’s representation is composed in the concept space.
+Also, when a new alignment goal is set, RLHF, SFT, and KE need to collect sufficient task samples and tune the LLM on the new dataset.
+In contrast, PaCE just needs to run the concept partitioner through CORD, which is expected to be much faster and more convenient.
+5.3 Context-dependent Concepts
+The structured activation space of LLMs and the large-scale concept dictionary of PaCE help to handle the influence of the context for a concept in the target prompt.
+As the LLM scales up, its capability to capture and utilize contextual information grows with the help of attention modules.
+The activation space, as already used for many representation manipulation methods, is expected to convey the underlying semantic information of concepts in the sentence (context).
+That is, the space hosting concept vectors is not collapsed, and it is structured to distinguish different concepts.
+The representation (activation) to be steered at inference time encodes the context and conveys the semantics of a concept based on the context.
+Then, since our overcomplete concept dictionary in PaCE widely covers concepts of various categories, the sparse coding on this dictionary will effectively analyze the target representation as the linear combination of these concepts.
+6 Conclusion
 In this paper, we present PaCE, an activation engineering framework designed for aligning LLMs by effectively and efficiently addressing undesirable representations while retaining linguistic capabilities.
@@
 Acknowledgments and Disclosure of Funding
-This research was supported by ARO MURI contract W911NF-17-1-0304, DARPA GARD HR001119S0026, DARPA RED HR00112090132, ODNI IARPA HIATUS Program Contract #2022-22072200005, and NSF Graduate Research Fellowship #DGE2139757.
-The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of NSF, ODNI, IARPA, or the U.S.
+This research was supported by ARO MURI W911NF-17-1-0304, DARPA GARD HR001119S0026, DARPA RED HR00112090132, ODNI IARPA HIATUS #2022-22072200005, the NSF grant 2031985, Simons Foundation MoDL 135615, a gift from AWS AI to Penn Engineering’s ASSET Center for Trustworthy AI, and NSF Graduate Research Fellowship #DGE2139757.
+We would like to thank Liangzu Peng, Hancheng Min, Bowen Li, Xinyu Yang, and Fengrui Tian for their suggestions in the presentation and experiments.
+The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, NSF, or the U.S.
 Government.
@@
 B.4 Implementation Details
-In our experiments, each response of the target LLM is set at a maximum of $512$ tokens.
+The experiments are conducted on a workstation of 8 NVIDIA A40 GPUs.
+Each response of the target LLM is set at a maximum of $512$ tokens.
+Activation vectors are extracted from the last-$29^{\text{th}}$ to the last-$11^{\text{th}}$ layer (totaling $19$ layers) of the target LLM’s decoder layers.
+All LLaMA-2 models in our experiments are the chat version (i.e., optimized for dialogue use cases).
+Concept Dictionary Construction and Partition.
 We set the scalar of the representation reading for concept vectors to $3.0$.
-The experiments are conducted on a workstation of 8 NVIDIA A40 GPUs.
-Activation vectors are extracted from the last $19$ layers of the target LLM’s decoder layer.
-For each input prompt, the decomposition is conducted on the inference process of the first next token, and the linear weights are reused for all next token predictions.
-All LLaMA-2 models in our experiments are the chat version (i.e., optimized for dialogue use cases).
 GPT-4-0125 is used for dictionary construction and concept partition.
-All alignment experiments use the top $10,000$ concepts from our PaCE-1M dataset to construct the concept dictionary (as Table 3 validates that the performance is high and does not change much after dictionary size $\geq 9000$).
 Each concept of PaCE-1M has at least 30 contextual sentences.
 For each alignment task, PaCE removes the top 50 undesirable concepts ranked by the GPT partitioner (§3).
-When solving the optimization problem for decomposition in §3.3, we set $\tau=0.95$ and $\alpha=0.05$ following the observations in [82].
-The MMLU evaluation is the 5-shot setting where 5 demonstrations are provided during question prompting.
 After retrieving the relevant knowledge (with the contriever [27]) from Wikipedia for concept synthesis, we take the top-5 ranked facts to append the instruction of LLM.
 The FAISS-indexed [31] Wikipedia is a snapshot of the 21 million disjoint text blocks from Wikipedia until December 2018.
+Activation Decomposition and Intervention.
+For each input prompt, the decomposition is conducted on the inference process of the first next token, and the linear weights are reused for all next token predictions.
+All alignment experiments use the top $10,000$ concepts from our PaCE-1M dataset to construct the concept dictionary (as Table 3 validates that the performance is high and does not change much after dictionary size $\geq 9000$).
+When solving the optimization problem for decomposition in §3.3, we set $\tau=0.95$ and $\alpha=0.05$ following the observations in [82].
+Evaluation.
+The MMLU evaluation is the 5-shot setting where 5 demonstrations are provided during question prompting.
 For the prompting baseline in Table 1 and Table 4, the instruction to the target LLM is to let the model be aware of the partitioned undesirable concepts and not to respond contents relevant to these concept:
@@
 Then the design of Decomposition on $10^{4}$ Concepts means that the dictionary is updated to be the top $10,000$ concepts in our PaCE-1M dataset and the concept ‘harmful’ from our dataset is removed.
-The Clustering of Concepts indicates that we run subspace clustering (detailed in Appendix C.2) and manually choose to remove all concepts of the cluster 125 with the PaCE-solved coefficients: ‘murder’, ‘evil’, ‘kill’, ‘violence’, ‘dirty’, ‘bomb’, ‘violent’, ‘armed’, ‘gross’, ‘savage’, ‘vicious’, ‘explosive’, ‘abuse’, ‘assault’, ‘penetration’, ‘cruelty’, ‘corruption’, ‘tyranny’, ‘tortured’, ‘notorious’, ‘militant’, ‘bloody’, ‘insult’, ‘lure’, ‘ruthless’, ‘inhuman’, and ’brutal’.
+The Clustering of Concepts indicates that we run subspace clustering (detailed in Appendix C.2) and manually choose to remove all concepts of the cluster 125 with the PaCE-solved coefficients: ‘murder’, ‘evil’, ‘kill’, ‘violence’, ‘dirty’, ‘bomb’, ‘violent’, ‘armed’, ‘gross’, ‘savage’, ‘vicious’, ‘explosive’, ‘abuse’, ‘assault’, ‘penetration’, ‘cruelty’, ‘corruption’, ‘tyranny’, ‘tortured’, ‘notorious’, ‘militant’, ‘bloody’, ‘insult’, ‘lure’, ‘ruthless’, ‘inhuman’, and ‘brutal’.
 Concept Partitioner means that we instruct GPT-4 to classify every concept as benign or undesirable (with a ranking score) and remove the top $10$ undesirable concepts with the PaCE-solved weights.
 Lastly, the Removal of Top $50$ Concepts suggests that we remove the top 50 concepts in the undesirable partition.
-Table 3 shows the effect of the dictionary size on three metrics (safety score, response fluency, and the average time per response).
+Figure 3 shows the effect of the dictionary size on three metrics (safety score, response fluency, and the average time per response).
 The fluency metric remains relatively consistent across different dictionary sizes, showing that PaCE’s decomposition maintains the general linguistic performance.
@@
 This validates our experiment choice of the dictionary size in this interval.
-B.6 Limitations, Societal Impacts, and Future Works
+Figure 11: Ablation study for the regularization $\tau$.
+$\tau$ Note Safety (%, $\uparrow$)
+0 Pure $\ell_{2}$ 68.9
+0.35 N.A. 65.4
+0.65 N.A. 71.6
+0.95 N.A. 72.0
+1.0 Pure $\ell_{1}$ 66.5
+Figure 11 shows that the regularization with $\tau=0.95$ yields the best safety performance among the five choices.
+Pure ridge regression ($\tau=0$) and pure lasso regression ($\tau=1$) do not perform as well as the mixed regularization strategy.
+B.6 Providing Dictionary and Scores to Target LLM
+For clarity, we elaborate on more details of how PaCE uses the concept dictionary.
+In our paper, dictionaries are a collection of concept vectors and are frozen for representation decomposition.
+First, the LLM takes an input prompt (e.g., malicious requests).
+Then an activation engineering framework [88, 72] extracts the activations at each decoder block of the transformer.
+Such extraction results in a vector corresponding to the input prompt, which can then be modified for steering in different ways.
+For PaCE, the steering has two main stages.
+Stage 1 pre-computes the large-scale concept dictionary offline and the partition (i.e., scores) of which concepts represent benign/harmful concepts.
+Stage 2 extracts the representation of an input prompt at inference time and uses sparse coding to decompose this as the linear combination of atoms in our frozen dictionary.
+We then modify this linear combination by removing undesirable components and proceeding with inference in the LLM with the detoxified representation.
+B.7 Limitations, Societal Impacts, and Future Works
 While our framework shows promising results, there exist potential limitations and several directions worth further exploration to address them.
@@
 In this paper, we follow the current practice (§2.2) to represent a concept by a single vector.
-Nonetheless, there might be several alternatives.
+Nonetheless, several alternatives could be explored.
 Results on linear polysemy [2, 15, 84] suggest that a concept might be better represented by multiple vectors or low-dimensional linear subspaces, each corresponding to different semantic meanings.
@@
 This is expected, since i) $\left\lVert{\boldsymbol{v}}_{i}^{\ell}\right\rVert_{2}=1$, so $\left\lVert{\boldsymbol{v}}^{\text{all}}_{i}\right\rVert_{2}=19$, ii) the linear dimensionality reduction preserves most of the energy.
-Algorithm. We apply Elastic Net Subspace Clustering (EnSC) on the preprocessed vectors to obtain $200$ clusters. The parameters of EnSC is set to $\tau=1$ and $\gamma=100$.
+Algorithm. We apply Elastic Net Subspace Clustering (EnSC) [82] on the preprocessed vectors to obtain $200$ clusters. The parameters of EnSC is set to $\tau=1$ and $\gamma=100$.
 Results.
@@
 This section presents the textual results generated using PaCE. It includes detailed detoxification comparisons with baseline models and analyses of the emergent clusters from the dataset.
-D.1 Baseline Responses
+D.1 Baseline Responses and Additional Benchmark
 Figure 13 shows the full response version of the Figure 5.
@@
 We observe that PaCE outperforms in detoxification performance by not outputting controversial terms, while maintaining general linguistic capabilities compared to other baselines.
+Table 5: Detoxification evaluation for PaCE, representation manipulation, and training-free baselines on AdvBench.
+Vanilla PE VecAdd OrthoProj PaCE (Ours)
+LlaMA2-7B-Chat 11.72 91.90 94.51 92.81 96.65
+LlaMA2-13B-Chat 18.04 93.86 95.33 96.72 99.17
+AdvBench [89] adversarially optimizes a jailbreak suffix for a harmful behavior request.
+Table 5 shows the LlaMA-7B-Chat and LlaMA-13B-Chat safety scores (%, $\uparrow$) on the effective set of suffix attacks for AdvBench harmful behavior set.
+The detoxification setup follows §4.1.
+We observe that PaCE outperforms other baselines.
+We also note that the outperformance of PaCE in §4.1’s jailbreaks is more significant than that in suffix attacks.
+This is potentially because story-telling and roleplay jailbreaks contain more complex and entangled concepts.
+Under this scenario, PaCE decomposes the target representation and well estimates the malicious component, while VecAdd and OrthoProj do not model the space sufficiently.
+In the AdvBench case, instead, the optimized adversarial suffix can be regarded as the text-space inversion of straightforward malicious concepts.
+PaCE and other defense mechanisms in latent space and prompt space shall effectively defend these suffixes more easily.
 D.2 Concept Clustering